OpenStudy (anonymous):

x - y = 4 x + y = 8 The x-coordinate of the solution to the system shown is _____.

OpenStudy (anonymous):

think of two numbers that add to 8 and are 4 apart probably can guess in one or at most two tries

OpenStudy (chosenmatt):

what is 8-4? then what is 4+4? then you will get your answer:)

OpenStudy (midhun.madhu1987):

Add both the equations and y will be eliminated. Find x.

OpenStudy (cuanchi):

y=8-x x-(8-x)=4 x-8+x=4 2x=4+8 x=(4+8)/2 x=6 y=8-x y=8-6=2 x=6 y=2

OpenStudy (midhun.madhu1987):